OpenStudy (anonymous):

Medal Solve. A pharmacist receives an IV order from a doctor for 100 mg of cefaclor. The label on the bottle of cefaclor has a concentration of 20 mg of drug per 250 mL of liquid. How many mL of the liquid should the pharmacist prepare for the IV? ?mL

OpenStudy (freckles):

what is 100/20?

OpenStudy (freckles):

basically you want to solve:\[\frac{100}{20}=\frac{x}{250}\]

OpenStudy (anonymous):

1250

OpenStudy (anonymous):

am i right

OpenStudy (freckles):

looks fine

OpenStudy (anonymous):

so the answer is 1250ml

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thank you

OpenStudy (jdoe0001):

\(\large { \begin{array}{ccllll} mg&mL \\\hline\\ 20&250\\ 100&x \end{array}\implies \cfrac{20}{100}=\cfrac{250}{x}\to 1250 }\) same thing pretty much

OpenStudy (anonymous):

so x is 1250 or 100

OpenStudy (jdoe0001):

x is 1250